OpenStudy (anonymous):

Can you explain this question: solve the equation by the square root property? 4(x-2)^2 =20.

OpenStudy (anonymous):

(x-2)^2 =20/4 \[X-2=\pm \sqrt{5}\] \[X=2\pm \sqrt{5}\]

OpenStudy (anonymous):

ty